OpenStudy (anonymous):

Find the derivative. F(x)=arcsine(radical x^2 +1) Can anybody find it?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

Let\[u =\sqrt{x ^{2}+1}\]

OpenStudy (anonymous):

Find derivative u. There is a formula for derivative of sine inverse.\[(d/dx)(\sin^{-1} u)=(1/\sqrt{1-u ^{2}})(du/dx) \]

OpenStudy (anonymous):

Yes, but i keep getting (1/ radical -1 - x^2) is that fine>

OpenStudy (anonymous):

* ?

OpenStudy (anonymous):

\[x/[\sqrt{(1-x ^{2}-1)}\sqrt{x ^{2}+1}]\]Simplify

OpenStudy (anonymous):

Oh, I see thank you.